if the population S(t) of a species of mosquitos in the Okavango region Delta at any time t >0 is modeled by the differential equation θ(ds/dt)+γS=〖αe〗^(-φt), whereθ,γ and α are positive constants and φ is a nonnegative constant, and if the initial population of this species is S(0)=S_0,
show that the mosquito population S(t) at any time t >0 is given by S(t)=α/(γ-θφ) e^(-φt)+(S_0-α/(γ-θφ) ) e^(-φt)

show that if φ=0 the species population S(t) approaches α/γ as t →∞ but if φ> the S(t) approaches zero as t→∞
1
Expert's answer
2020-12-01T03:06:50-0500

Let us consider the differential equation


"\Theta\frac{dS}{dt}+\gamma S=\alpha e^{-\varphi t}", "S(0)=S_0"


where "\varphi\ge 0,\ \Theta>0,\ \gamma>0,\ \alpha >0."


Let us multiply the both part of equation by "\frac{1}{\Theta}e^{\frac{\gamma t}{\Theta}}:"


"\frac{dS}{dt}e^{\frac{\gamma t}{\Theta}}+\frac{\gamma}{\Theta} Se^{\frac{\gamma t}{\Theta}}=\frac{\alpha}{\Theta} e^{-\varphi t}e^{\frac{\gamma t}{\Theta}}".


The last equation is equivalent to


"\frac{d}{dt}(Se^{\frac{\gamma t}{\Theta}})=\frac{\alpha}{\Theta} e^{(\frac{\gamma }{\Theta}-\varphi) t}"


Therefore,


"Se^{\frac{\gamma t}{\Theta}}=\frac{\alpha}{\gamma -\varphi \Theta} e^{\frac{\gamma-\varphi\Theta }{\Theta} t}+C"


Let us multiply the both part of equation by "e^{-\frac{\gamma t}{\Theta}}:"


"S(t)=\frac{\alpha}{\gamma -\varphi \Theta} e^{-\varphi t}+Ce^{-\frac{\gamma t}{\Theta}}".


Let "t=0". Then


"S_0=S(0)=\frac{\alpha}{\gamma -\varphi \Theta} +C," and thus "C=S_0-\frac{\alpha}{\gamma -\varphi \Theta}".


Consequently, the solution of the differential equation is the following:


"S(t)=\frac{\alpha}{\gamma -\varphi \Theta} e^{-\varphi t}+(S_0-\frac{\alpha}{\gamma -\varphi \Theta})e^{-\frac{\gamma t}{\Theta}}".


If "\varphi =0", then "S(t)=\frac{\alpha}{\gamma} +(S_0-\frac{\alpha}{\gamma})e^{-\frac{\gamma t}{\Theta}}", and thus


"\lim_{t\to+\infty}S(t)=\lim_{t\to+\infty}[\frac{\alpha}{\gamma} +(S_0-\frac{\alpha}{\gamma})e^{-\frac{\gamma t}{\Theta}}]=\frac{\alpha}{\gamma}"


So, if "\u03c6=0" the species population "S(t)" approaches "\frac{\alpha}{\gamma}" as "t\to+\infty."



If "\varphi>0," then "\lim_{t\to+\infty}S(t)=\lim_{t\to+\infty}[\frac{\alpha}{\gamma -\varphi \Theta} e^{-\varphi t}+(S_0-\frac{\alpha}{\gamma -\varphi \Theta})e^{-\frac{\gamma t}{\Theta}}]=0".


Therefore, if "\u03c6>0" the "S(t)" approaches zero as "t\to+\infty."